40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 29 19:38:09 2015, gautam, Update, General, FOL fiber box revamp IMG_0017.JPGFOL_schematic.pdfIMG_0018.JPGIMG_0016.JPG
    Reply  Wed Sep 30 13:07:13 2015, gautam, Update, General, FOL fiber box revamp FOL_schematic.pdf
       Reply  Tue Oct 6 16:56:40 2015, gautam, Update, General, FOL fiber box revamp 20151006_Xbeat_psd.pdf20151006_Xbeat_coherence.pdf
Message ID: 11650     Entry time: Tue Sep 29 19:38:09 2015     Reply to this: 11652
Author: gautam 
Type: Update 
Category: General 
Subject: FOL fiber box revamp 

The new 2x2 fiber couplers arrived today so Eric gave me an overview on the changes to be made to the existing configuration of the FOL fiber box. I removed the box from the table after ensuring that the PDs were powered OFF and removing and capping all fiber leads on the front panel. Here is a summary of the changes made.

  • On-Off positions for the rocker switches corrected - these switches for the power to the PDs were installed such that the "1" position was OFF. I flipped both the switches such that the "1" position now corresponds to ON (see Attachment #1).
  • All the couplers/beam combiners/splitters were initially removed. 
  • I then re-configured the layout as per the schematic (Attachment #2). I only needed to use one of the 4 new 2x2 couplers ordered. I think the 1x2 couplers are appropriate for mixing the PSL and AUX beams, as if we use a 2x2 coupler, half of the mixed light goes nowhere? Indeed, if we had one more such coupler, we could do away with the 2x2 coupler I am now using to divide the PSL light into two. 
  • The spec-sheets on the inside of the top cover were updated to reflect the new hardware (Attachment #3).
  • The old hardware from the box that was not used, along with their spec-sheets, are stored temporarily in a Thorlabs lab snacks box (all the fibers have been capped).
  • The finished layout is shown in Attachment #4.

I then ran a quick check to see what the power levels were at the input to the PDs, using the fiber coupled power meter. However, I found that there was no light in the fiber marked "PSL light in" (the power meter read out "Sig. Low"). The X arm Aux light had an input power of 1.12 mW, which after the various coupling losses etc went down to 63 uW just before the PD. The corresponding figures for the Y arm are 200 uW and 2.2 uW. I am not too sure of how the AUX light is coupled into fibers so I am not trying to tweak the alignment to see if I get more power. 

Attachment 1: IMG_0017.JPG  77 kB  | Hide | Hide all
IMG_0017.JPG
Attachment 2: FOL_schematic.pdf  56 kB  | Hide | Hide all
FOL_schematic.pdf
Attachment 3: IMG_0018.JPG  90 kB  | Hide | Hide all
IMG_0018.JPG
Attachment 4: IMG_0016.JPG  146 kB  | Hide | Hide all
IMG_0016.JPG
ELOG V3.1.3-